OpenStudy (anonymous):

The equation 3^(x-1) + 5^(x-1) = 34 has how many sol^n?

OpenStudy (anonymous):

if \(x>3\) then \(3^{x-1}+5^{x-1} >34 \) if \(x<3\) then \(3^{x-1}+5^{x-1} <34 \) so clearly \(x=3\)

mathslover (mathslover):

The answer given by @mukushla is verified correctly : http://www.wolframalpha.com/input/?i=solve+%3A+3%5E%7Bx-1%7D%2B5%5E%7Bx-1%7D%3D34+

OpenStudy (anonymous):

@mukushla i nt understand

OpenStudy (anonymous):

see if \(x>3\) then \(3^{x-1}+5^{x-1}>3^{2}+5^{2}=9+25=34 \) since \( x-1>2\) and the same thing for \(x<3 \)

OpenStudy (anonymous):

Why did u take it as 3

OpenStudy (anonymous):

because i knew that ... \(3^2+5^2=34 \)

OpenStudy (anonymous):

thxx

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

so.........hw and sol^n

OpenStudy (anonymous):

3...?

OpenStudy (anonymous):

yes just x=3

OpenStudy (anonymous):

@mukushla u r frm India

OpenStudy (anonymous):

nope...im from iran